# Basis Selection: Low-Rank Decomposition of Pretrained Large Language Models for Target Applications

### v2

@@
11footnotetext: Equal contribution.22footnotetext: Corresponding author. Email: yangli1@iastate.edu
@@
Iowa State University / Meta
@@
Large language models (LLMs) significantly enhance the performance of various applications, but they are computationally intensive and energy-demanding. This makes it challenging to deploy them on devices with limited resources, such as personal computers and mobile/wearable devices, and results in substantial inference costs in resource-rich environments like cloud servers. To extend the use of LLMs, we introduce a low-rank decomposition approach to effectively compress these models, tailored to the requirements of specific applications. We observe that LLMs pretrained on general datasets contain many redundant components not needed for particular applications. Our method focuses on identifying and removing these redundant parts, retaining only the necessary elements for the target applications. Specifically, we represent the weight matrices of LLMs as a linear combination of base components. We then prune the irrelevant bases and enhance the model with new bases beneficial for specific applications. Deep compression results on the Llama 2-7b and -13B models, conducted on target applications including mathematical reasoning and code generation, show that our method significantly reduces model size while maintaining comparable accuracy to state-of-the-art low-rank compression techniques.
@@ 1 Introduction
Large language models (LLMs) have significantly enhanced the performance of various applications in natural language processing, computer vision, and beyond. However, their large model sizes pose a bottleneck for many practical uses. The substantial computing resources required for LLM inference make it challenging to deploy them on devices with limited capabilities, such as personal computers and mobile/wearable devices. Moreover, even on hardware platforms with ample computing power, deploying LLMs consumes a significant amount of energy, raising concerns about sustainability. Therefore, it is essential to reduce the size of LLMs after pretraining to ease their computational demands and lower energy consumption.
Our approach exploits the relationship between pretrained models and specific target applications. Large language models (LLMs) are typically pretrained on vast datasets encompassing a wide range of tasks, many of which share common characteristics. This shared pretraining fosters synergies that enhance the performance of LLMs. However, the diversity among these tasks also introduces redundancy into the models. As demonstrated by our interpretation results in Section 3 shows, LLMs contain a significant number of redundant components that are unnecessary for a specific target application. By removing these redundant parts and retaining only the relevant ones, we can reduce the model’s size while preserving its performance on the target application. Since many scenarios only require support for a specific type of application, this approach effectively lowers the computing resource requirements and reduces inference costs.
@@
However, how do we identify the beneficial and redundant components of LLMs for a specific application? In this work, we address this problem through the lens of matrix factorization. Singular Value Decomposition (SVD) (Golub & Van Loan, 1996) factorizes a weight matrix $\mathbf{W}$ into the product of three matrices $\mathbf{U}$, $\mathbf{S}$, and $\mathbf{V}$, i.e., $\mathbf{W}=\mathbf{USV}^{\top}=\sum_{i}{s_{i}\mathbf{u}_{i}{\mathbf{v}_{i}^{\top}}}$, where $s_{i}$ are (positive) singular values, and $\mathbf{u_{i}}$ and $\mathbf{v_{i}}$ are column vectors of $\mathbf{U}$ and $\mathbf{V}$ with unit norms. Our interpretation results show that these column vectors $\mathbf{u_{i}}$ and $\mathbf{v_{i}}$ may carry specific meanings. For instance, in the LLaMA 2-7B model (Touvron et al., 2023b), when factorizing the weight matrix $\mathbf{W_{O}^{h}W_{V}^{h}}$ of an attention head that is likely useful for code generation tasks, de-embedding the resulting column vectors $\mathbf{u_{i}}$ and $\mathbf{v_{i}}$ reveals tokens like _in, <0x0A>, _and, _to, and _for. These column vectors are evidently highly useful for code generation but may be less relevant for tasks such as mathematical reasoning.
@@
We evaluate Basel on two models—Llama 2-7B and Llama 2-13B (Touvron et al., 2023b)—and two tasks—mathematical reasoning and code generation. We evaluate the pass@1 accuracy of the compressed models on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) for the mathematical reasoning task and HumanEval (Chen et al., 2021a) and MBPP (Austin et al., 2021) for the code generation task. Compared to state-of-the-art baselines, our approach achieves substantially better performance, improving accuracy by up to 19% when the compression ratio [^1] The compression ratio is defined as the ratio of the original model size to the compressed model size. exceeds 6 for mathematical reasoning and 4 for code generation. This also indicates that, in cases of deep compression, our method reduces model size by up to 2.7 times while maintaining comparable accuracy to baseline methods.
@@ 2 Related Work
Singular value decomposition (SVD) (Golub & Van Loan, 1996) has been applied to reduce the size of machine learning models. Prior research (Xue et al., 2013; Jaderberg et al., 2014; Denton et al., 2014; Zhang et al., 2015; Povey et al., 2018; Chen et al., 2018; Acharya et al., 2019; Noach & Goldberg, 2020) has developed various SVD algorithms to compress different components of models, such as DNNs, CNNs, and embedding layers for a range of applications including natural language processing, speech, and vision. The primary distinction between our work and these prior studies is that they do not relearn the importance of bases using the training data of target applications. Instead, they typically prune bases according to the singular values from the original or finetuned models. FWSVD (Hsu et al., 2022) evaluates the importance of individual weights rather than bases during SVD. As highlighted in Sections 1 and 3, the bases hold significant physical meanings. Considering importance at this level of granularity results in improved performance. As shown in Section 4, our approach surpasses FWSVD in deep compression performance.
A recent study (Sharma et al., 2024) applied SVD to large language models. Its focus is on determining the optimal rank for each layer, while our emphasis is on basis selection. The two methods are orthogonal but complementary and can be combined. Chen et al. (2021b) and Yu & Wu (2023) suggest reconstructing bases based on feature mimicking. These approaches are orthogonal to ours—they concentrate on basis reconstruction, whereas we focus on basis selection. Their methods complement ours and can be integrated together to achieve enhanced compression results.
@@ 3 Basel
where $\mathbf{U}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $\mathbf{S}=\mathrm{diag}\left(s_{1},\cdots,s_{r}\right)$, and $\mathbf{V}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$. The values $\left\{s_{i}\in\mathbb{R},i=1,\cdots,r\right\}$ are positive singular values. [^2] We drop zero singular values and the corresponding columns of matrices $U$ and $V$. The vectors $\left\{\mathbf{u_{i}}\in\mathbb{R}^{n},i=1,\cdots,r\right\}$ and $\left\{\mathbf{v_{i}}\in\mathbb{R}^{m},i=1,\cdots,r\right\}$ are orthonormal, i.e., $\left\|\mathbf{u_{i}}\right\|=1$, $\left\|\mathbf{v_{i}}\right\|=1$, $\mathbf{u_{i}}\perp\mathbf{u_{j}}$, and $\mathbf{v_{i}}\perp\mathbf{v_{j}}$ if $i\neq j$.
@@
Let matrix $\mathbf{W_{i}}=\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$, then
$\left\|\mathbf{W_{i}}\right\|=\sqrt{\mathrm{tr}\left(\mathbf{W_{i}}^{\top}\mathbf{W_{i}}\right)}=\sqrt{\mathrm{tr}\left(\mathbf{v_{i}}\mathbf{u_{i}}^{\top}\mathbf{u_{i}}\mathbf{v_{i}}^{\top}\right)}=\sqrt{\mathrm{tr}(\mathbf{v}_{i}\mathbf{v_{i}}^{\mathrm{T}})}=\sqrt{\mathrm{tr}(\mathbf{v_{i}}^{\mathrm{T}}\mathbf{v_{i}})}=1$
$\left\langle\mathbf{W_{i},W_{j}}\right\rangle=\mathrm{tr}\left(\mathbf{W_{i}}^{\top}\mathbf{W_{j}}\right)=\mathrm{tr}\left(\mathbf{v_{i}}\mathbf{u_{i}}^{\top}\mathbf{u_{j}}\mathbf{v_{j}}^{\top}\right)=0,\quad\text{if}\;i\neq j$
@@
Large language models pretrained on diverse datasets learn bases that capture a wide range of semantic concepts. To examine this, we factorize the $\mathbf{W_{O}^{h}W_{V}^{h}}$ matrix from the attention layers of both the original Llama 2-7B model and its math-finetuned counterpart. Focusing on bases associated with large singular values, we extract the corresponding $\mathbf{u}$ vectors and interpret their semantics via a de-embedding process. Specifically, we project each $\mathbf{u}$ vector into the vocabulary space by passing it through the output layer (i.e., the lm_head), yielding a logit vector over the vocabulary. We then identify the token with the highest logit as the de-embedding result. Table 1 summarizes our findings. First, some bases encode highly specific semantic content, such as concepts related to Apple, Python, programming, locations, non-English characters, and math symbols. Second, even within the same layer, different bases can represent very different meanings. For example, in the math-finetuned model at layer 22, attention head 6, basis 1 corresponds to non-English characters, while basis 3 corresponds to math symbols. Third, fine-tuning does not necessarily remove irrelevant bases: the non-English character basis remains present even after fine-tuning on math-focused data.
@@
Apple
Vanilla model, Layer 16,
Head 25, Basis 2
_iOS_Xcode_ios_Apple_Mac _iPhone_app_xcode_NS_App
Python
Vanilla model, Layer 17,
Head 6, Basis 1
._in<0x0A>_…_andĽ_tofor!
Location
Vanilla model, Layer 17,
Head 25, Basis 1
_Massachusetts_Illinois_Chicago_Boston_Dan _Harvard_Connecticut_IL_Bulg_Bulgar
Non-English
Math-finetuned model,
Layer 22, Head 6, Basis 1
:Yā’
Math
Math-finetuned model,
Layer 22, Head 6, Basis 3
@@
These findings suggest that many bases in the model are useful for specific tasks, but may be irrelevant for others. When these irrelevant bases are used as filters in non-target applications, two scenarios can occur: the filter may not be activated (due to a small inner product $\left\langle\mathbf{x},\mathbf{v}_{i}\right\rangle$), or worse, the filter is activated, introducing harmful information into the output and degrading performance. This indicates that pruning such bases could reduce model size with minimal performance loss, and in some cases, even enhance performance for the target application.
@@
$\mathbf{\widetilde{W}}=\sum_{i=1}^{r}\tilde{{s}}_{i}\mathbf{u_{i}}\mathbf{v_{i}}^{\top}+\sum_{j=1}^{\tilde{r}}\tilde{\mathbf{u}}_{\mathbf{{j}}}\tilde{\mathbf{v}}_{\mathbf{{j}}}^{\top}$ (3)
@@
Data: KeepRatio, PruningTimes, KeepingEpoch, PruningEpoch, PostFineTuningEpoch
@@
4 Convert the weight matrix of each linear layer in MM—excluding the embedding layer—into the form specified by equation equation 3;
@@
6  Tune the learnable parameters in equation equation 3;
7
@@ 11  Tune the learnable parameters;
21
22for each layer do
23  Compute the low rank matrix $\widetilde{W}$ based on equation equation 3;
@@ 25  Use two linear layers to substitute for the original layer;
26  The first layer’s weight matrix is $S^{\prime}V^{\prime\textrm{T}}$;
27  The second layer’s weight matrix is $U^{\prime}$;
28
@@ 32  FineTune the new model $M^{\prime}$;
In the first term, $\mathbf{u_{i}}\mathbf{v_{i}}^{\mathrm{T}}$ represents the original bases in the pretrained model. To assess their importance, we initialize their weights $\tilde{s}_{i}$ with their original singular values and then retrain these weights (while keeping the bases fixed) on the training set of the target application. The aim is that, after retraining, the bases important for the target application will have larger singular values, whereas those that are useless or detrimental will have zero or very small singular values. This allows us to identify and prune the less useful bases. Relearning the importance of bases for the target application distinguishes our approach from previous methods. Prior approaches either use the singular values in the original model (Xue et al., 2013; Jaderberg et al., 2014; Denton et al., 2014; Zhang et al., 2015; Povey et al., 2018; Chen et al., 2018; Acharya et al., 2019; Noach & Goldberg, 2020; Sharma et al., 2024) or assess the importance of weight parameters, other than the importance of the bases, to prune them (Hsu et al., 2022). They do not relearn the importance of the bases specifically for the target application. From a signal processing perspective, this first term allows us to adjust the weight for each filter, catering to the needs of the target application.
Algorithm 1 presents our proposed approach. Basel operates on a pretrained or fine-tuned model as input. Similar to Hsu et al. (2022); Sharma et al. (2024), Basel iteratively prunes the bases of all linear layers—excluding the embedding layer—including the $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, $\mathbf{W_{V}}$, and $\mathbf{W_{O}}$ matrices in the attention layers, the linear layers in the feedforward layers, and the output layer. The same compression process is applied across these components. After each pruning step, the learnable parameters $\tilde{s}_{i}$, $\tilde{\mathbf{u}}_{j}$, and $\tilde{\mathbf{v}}_{j}$ are finetuned to compensate for performance degradation. Ultimately, a new weight matrix $\widetilde{\mathbf{W}}$ with a smaller rank $r^{\prime}$ is learned. We perform a standard SVD on it, representing it as the product of matrices $\mathbf{U^{\prime}}$, $\mathbf{S^{\prime}}$, and $\mathbf{V^{\prime}}^{\top}$. We then replace the original layer with two new layers: $\mathbf{S^{\prime}V^{\prime}}^{\top}$ becomes the weight matrix of the first new layer, and $\mathbf{U^{\prime}}$ becomes the weight matrix of the second new layer. This reduces the number of parameters from $nm$ to $(n+m)r^{\prime}$. The new model is subsequently further finetuned to enhance its performance on the target application.

### v4

@@
11footnotetext: Equal contribution.22footnotetext: Corresponding author. Address: 2434 Osborn Dr, Ames, IA 50011, United States. Email: yangli1@iastate.edu.
@@
Iowa State University and Meta
@@
As the size of language models increases, they deliver substantial performance improvements across a variety of applications. However, this growth also leads to greater computational demands, making deployment on resource-constrained devices—such as personal computers and mobile or wearable devices—more challenging, and significantly raising inference costs on cloud servers. To address these challenges, we introduce Basel, a method to streamline language models by leveraging the semantic structure of their weight matrices. Specifically, Basel treats each weight matrix as a linear combination of bases, selectively retaining those that are associated with essential semantics for the target application, pruning redundant ones, and introducing new bases that enhance task performance. Experimental results demonstrate that Basel achieves significant model size reduction compared to baseline techniques, while maintaining comparable or even superior accuracy across diverse applications.
@@ 1 Introduction
Large language models (LLMs) (llama; gpt4; gemini) have significantly enhanced the performance of various applications in natural language processing, computer vision, and beyond. However, their large model sizes pose a bottleneck for many practical uses. The substantial computing resources required for LLM inference make it challenging to deploy them on devices with limited capabilities, such as personal computers and mobile/wearable devices (Li2023FoldingAttention; li-etal-2025-breaking). Moreover, even on hardware platforms with ample computing power, deploying LLMs consumes a significant amount of energy, raising concerns about sustainability. Therefore, it is essential to reduce the size of LLMs after pretraining to ease their computational demands and lower energy consumption.
Our approach exploits the relationship between pretrained models and specific target applications. Large language models are typically pretrained on vast datasets encompassing a wide range of tasks, many of which share common characteristics. This shared pretraining fosters synergies that enhance the performance of LLMs. However, the diversity among these tasks also introduces redundancy into the models. As demonstrated by our results in Section 3, LLMs contain redundant components that are unnecessary for specific target applications. By removing these redundant parts and retaining only the relevant ones, we can reduce the model’s size while preserving its performance on the target application. Since many scenarios only require support for a specific type of application (chang-etal-2024-target), this approach effectively lowers the computing resource requirements and reduces inference costs.
@@
However, how do we identify the beneficial and redundant components of LLMs for a specific application? In this work, we address this problem through the lens of matrix factorization. Singular value decomposition (SVD) (SVD_algorithm) factorizes a weight matrix $\mathbf{W}$ into the product of three matrices $\mathbf{U}$, $\mathbf{S}$, and $\mathbf{V}$, i.e., $\mathbf{W}=\mathbf{USV}^{\top}=\sum_{i}{s_{i}\mathbf{u}_{i}{\mathbf{v}_{i}^{\top}}}$, where $s_{i}$ are (positive) singular values, and $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ are column vectors of $\mathbf{U}$ and $\mathbf{V}$ with unit norms. Our results show that these column vectors $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ may carry specific meanings. For instance, in the LLama 2-7B model (llama2), when factorizing the weight matrix $\mathbf{W}_{O}^{h}\mathbf{W}_{V}^{h}$ of an attention head that is likely useful for code generation tasks, de-embedding the resulting column vectors $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ reveals tokens like _in, <0x0A>, _and, _to, and _for. These column vectors are evidently useful for code generation but may be less relevant for tasks such as mathematical reasoning.
@@
[…]
@@
The final version of this work is published at Transactions on Machine Learning Research (TMLR) (basis_selection), and the source code of the work is publicly available at https://github.com/Iowa-State-University-AI-System-Group/Basel .
@@ 2 Related Work
Singular value decomposition (SVD) (SVD_algorithm) has been applied to reduce the size of machine learning models. Prior research (SVD0; SVD1; SVD2; SVD3; SVD4; SVD_recent1; SVD_recent2; SVD_recent3) has developed various SVD algorithms to compress different components of models, such as DNNs, CNNs, and embedding layers for a range of applications including natural language processing, speech, and vision. The primary distinction between our work and these prior studies is that they do not relearn the importance of bases using the training data of target applications. Instead, they typically prune bases according to the singular values from the original or finetuned models. FWSVD (FWSVD) evaluates the importance of individual weights rather than bases during SVD. Considering the importance of bases may improve performance. As shown in Section 4, our approach surpasses FWSVD in deep compression performance.
A recent study (LASER) applied SVD to large language models. Its focus is on determining the optimal rank for each layer, while our emphasis is on basis selection. The two methods are orthogonal but complementary and can be combined. Drone; AFM; asvd; svdllm; impact propose reconstructing bases by minimizing the discrepancy between activations before and after compression. These methods are also orthogonal to ours: they concentrate on basis reconstruction, whereas we focus on basis selection. In principle, their techniques may be integrated with ours to achieve stronger compression performance.
Beyond pruning bases, compression can also be achieved by pruning at other granularities, such as weight pruning (FLAP; wanda) and layer pruning (trimllm).
@@ 3 Basel
where $\mathbf{U}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $\mathbf{S}=\mathrm{diag}\left(s_{1},\cdots,s_{r}\right)$, and $\mathbf{V}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$. The values $\left\{s_{i}\in\mathbb{R},i=1,\cdots,r\right\}$ are positive singular values. [^1] We drop zero singular values and the corresponding columns of matrices $U$ and $V$. The vectors $\left\{\mathbf{u}_{i}\in\mathbb{R}^{n},i=1,\cdots,r\right\}$ and $\left\{\mathbf{v}_{i}\in\mathbb{R}^{m},i=1,\cdots,r\right\}$ are orthonormal, i.e., $L^{2}$ norms $\left\|\mathbf{u}_{i}\right\|_{2}=1$, $\left\|\mathbf{v}_{i}\right\|_{2}=1$, Euclidean inner products $\langle\mathbf{u}_{i},\mathbf{u}_{j}\rangle=\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle=0$, for $i\neq j$.
@@
Let matrix $\mathbf{W}_{i}=\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$, then
$\textrm{Frobenius norm}\left\|\mathbf{W}_{i}\right\|_{F}=\sqrt{\mathrm{tr}\left(\mathbf{W}_{i}^{\top}\mathbf{W}_{i}\right)}=\sqrt{\mathrm{tr}\left(\mathbf{v}_{i}\mathbf{u}_{i}^{\top}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}\right)}=\sqrt{\mathrm{tr}(\mathbf{v}_{i}\mathbf{v}_{i}^{\top})}=\sqrt{\mathrm{tr}(\mathbf{v}_{i}^{\top}\mathbf{v}_{i})}=1$
$\textrm{Frobenius inner product}\left\langle\mathbf{W}_{i},\mathbf{W}_{j}\right\rangle_{F}=\mathrm{tr}\left(\mathbf{W}_{i}^{\top}\mathbf{W}_{j}\right)=\mathrm{tr}\left(\mathbf{v}_{i}\mathbf{u}_{i}^{\top}\mathbf{u}_{j}\mathbf{v}_{j}^{\top}\right)=0,\quad\text{if}\;i\neq j$
@@
Large language models pretrained on diverse datasets learn bases that capture a wide range of semantic concepts. To examine this, we factorize the $\mathbf{W}_{O}^{h}\mathbf{W}_{V}^{h}$ matrix from the attention layers of both the original Llama 2-7B model and its math-finetuned counterpart. Focusing on bases associated with large singular values, we extract the corresponding $\mathbf{u}$ vectors and interpret their semantics via a de-embedding process. Specifically, we project each $\mathbf{u}$ vector into the vocabulary space by passing it through the output layer (i.e., the lm_head), yielding a logit vector over the vocabulary. We then identify the token with the highest logit as the de-embedding result. Table 1 summarizes our findings. First, some bases are associated with semantic content, such as concepts related to Python, locations, non-English characters, and math symbols. Second, even within the same layer, different bases may correspond to different meanings. For example, in the math-finetuned model at layer 22, attention head 6, basis 1 corresponds to non-English characters, while basis 3 corresponds to math symbols. Revealing the full semantic structure of these bases requires more in-depth analysis, such as the approaches explored in (oikarinen2025evaluating), which we leave for future work.
@@
Python Vanilla model, Layer 17, Head 6, Basis 1 ._in<0x0A>_…_andĽ_tofor!
Location Vanilla model, Layer 17, Head 25, Basis 1 _Massachusetts_Illinois_Chicago_Boston_Dan _Harvard_Connecticut_IL_Bulg_Bulgar
Non-English Math-finetuned model, Layer 22, Head 6, Basis 1 :Yā’
@@
These findings suggest that some bases in the model are useful for some tasks, but may be irrelevant for others. When these irrelevant bases are used as filters in non-target applications, two scenarios can occur: the filter may not be activated (due to a small inner product $\left\langle\mathbf{x},\mathbf{v}_{i}\right\rangle$), or worse, the filter is activated, introducing harmful information into the output and degrading performance. This indicates that pruning such bases could reduce model size with minimal performance loss, and in some cases, even enhance performance for the target application.
@@
$\mathbf{\widetilde{W}}=\sum_{i=1}^{r}\tilde{{s}}_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}+\sum_{j=1}^{\tilde{r}}\tilde{\mathbf{u}}_{j}\tilde{\mathbf{v}}_{j}^{\top}$ (3)
In the first term, $\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ represents the original bases in the pretrained model. To assess their importance, we initialize their weights $\tilde{s}_{i}$ with their original singular values and then retrain these weights (while keeping the bases fixed) on the training set of the target application. The aim is that, after retraining, the bases important for the target application will have larger singular values, whereas those that are useless or detrimental will have zero or very small singular values. This allows us to identify and prune the less useful bases. Relearning the importance of bases for the target application distinguishes our approach from previous methods. Prior approaches either use the singular values in the original model (SVD0; SVD1; SVD2; SVD3; SVD4; SVD_recent1; SVD_recent2; SVD_recent3; LASER) or assess the importance of weight parameters, other than the importance of the bases, to prune them (FWSVD). They do not relearn the importance of the bases useful for the target application. From a signal processing perspective, this first term allows us to adjust the weight for each filter, catering to the needs of the target application.
@@
Data: Hyperparameters including KeepRatio, PruningTimes, KeepingEpoch, PruningEpoch, PostFineTuningEpoch, $\tilde{r}$
@@
4 Convert the weight matrix of each linear layer in MM—excluding the embedding layer—into the form specified by equation (3);
@@
6  Tune the learnable parameters in equation (3) including $\tilde{s}_{i},\tilde{\mathbf{u}}_{j},\tilde{\mathbf{v}}_{j},\;i=1,\cdots,r,\;j=1,\cdots,\tilde{r}$;
@@ 10  Tune the learnable parameters;
20
21for each layer do
22  Compute the low rank matrix $\mathbf{\widetilde{W}}$ based on equation (3);
@@ 24  Use two linear layers to substitute for the original layer;
25  The first layer’s weight matrix is $\mathbf{S^{\prime}V^{\prime}}^{\top}$;
26  The second layer’s weight matrix is $\mathbf{U^{\prime}}$;
27
@@ 31  FineTune the new model $M^{\prime}$;
Algorithm 1 presents our proposed approach. Basel operates on a pretrained or fine-tuned model as input. Similar to FWSVD; LASER, Basel iteratively prunes the bases of all linear layers—excluding the embedding layer—including the $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$, and $\mathbf{W}_{O}$ matrices in the attention layers [^2] In our implementation, matrices $\mathbf{W}_{V}$ and $\mathbf{W}_{O}$ are pruned independently. Exploring alternative strategies, such as jointly pruning $\mathbf{W}_{V}$ and $\mathbf{W}_{O}$, is left for future work., the linear layers in the feedforward layers, and the output layer. The same compression process is applied across these components. After each pruning step, the learnable parameters $\tilde{s}_{i}$, $\tilde{\mathbf{u}}_{j}$, and $\tilde{\mathbf{v}}_{j}$ are finetuned to compensate for performance degradation. Ultimately, a new weight matrix $\widetilde{\mathbf{W}}$ with a smaller rank $r^{\prime}$ is learned. We perform a standard SVD on it, representing it as the product of matrices $\mathbf{U^{\prime}}$, $\mathbf{S^{\prime}}$, and $\mathbf{V^{\prime}}^{\top}$. We then replace the original layer with two new layers: $\mathbf{S^{\prime}V^{\prime}}^{\top}$ becomes the weight matrix of the first new layer, and $\mathbf{U^{\prime}}$ becomes the weight matrix of the second new layer. This reduces the number of parameters from $nm$ to $(n+m)r^{\prime}$. The new model is subsequently further finetuned to enhance its performance on the target application.
Overhead Analysis. Basel updates only the singular values and the newly introduced bases, while keeping the original bases frozen. Unlike full fine-tuning, Basel involves far fewer learnable parameters, substantially reducing training overhead. Table 2 compares Basel and full fine-tuning on Llama-2-7B in terms of GPU hours and memory consumption, evaluated on NVIDIA L40S GPUs (batch size = 32, max sequence length = 512). Under this setting, Basel fits on a single L40S GPU, whereas full fine-tuning requires at least three L40S GPUs due to its higher memory demand. Overall, Basel consumes only about 46% of the GPU hours and 30% of the GPU memory used by full fine-tuning.
Таблица 2: GPU hours and GPU memory consumption of Basel versus full fine-tuning on Llama 2-7B using NVIDIA L40S GPUs (batch size = 32, max sequence length = 512).
GPU Hours per Batch Total GPU Memory (GB)
Basel $2.87\times 10^{-3}$ 40.7
Full fine-tuning $6.19\times 10^{-3}$ 136.2